OpenStudy (anonymous):

solve for y. x=yt+25

OpenStudy (anonymous):

Why don't you try the first step. Get yt on one side by itself

OpenStudy (anonymous):

y=x- 25/t

OpenStudy (anonymous):

im sorry its y=yt+2s

OpenStudy (anonymous):

I know what you are saying Ankur but you need to use brackets to make it clear. It doesn't really help zaynab to just jump to the end though

OpenStudy (anonymous):

yes...then replace 25 with 2s..simple

OpenStudy (anonymous):

ok...thanx chaguanas...

OpenStudy (anonymous):

let me rewrite it again

OpenStudy (anonymous):

y=(x- 2s)/ t

OpenStudy (anonymous):

smone fan me plsssssss

OpenStudy (anonymous):

thanxxxxxxxxxxxxxxxxxx

OpenStudy (anonymous):

u welcome and thanks for answering my question

OpenStudy (anonymous):

no prob...

OpenStudy (anonymous):

i think v should fan each other to keep the morale high